Evaluate the following integral:

∫ √9x−5dx.

 


a.    1 / 6 √(9x−5) + c


b.   1 / 6 √(9x−5)^3 + c


c.    2 / 27 √ (9x−5) + c


d.   2 / 27 √ (9x - 5) ^3 + c



1
Expert's answer
2022-05-01T17:08:58-0400

"F(x)=\sqrt{9x-5}"

"\int\sqrt{9x-5} dx=2/27(\sqrt{(9x-5)^3}+C"

u=9x-5

dx=1/9 u

"1/9\smallint \sqrt{u}du=2/27 \sqrt{u^3}=2/27 \sqrt{(9x-5)^3} +C"


Answer: d